# ?reference?

### v1

@@ 5 Experimental Setup
Extrapolated-Flowchart Extrapolated-Results Extrapolated-Architecture Summary
@@
Model Extrapolated-Flowchart Extrapolated-Summary Extrapolated-Architecture Extrapolated-Results
@@ C F L C F L C F L C F L
For generating first-draft of diagrams, we experiment with state-of-the-art models to extract table from existing figures during diagram planning (Figure 16): 1) MatCha Liu et al. (2023b), 2) Chart-to-Text Kantharaj et al. (2022), 3) UniChart Masry et al. (2023), 4) DePlot Liu et al. (2023a) and GPT4-V OpenAI (2023). Finally, based on our assessment (Figure 18), we concluded that GPT4-V outperformed other models and was the best choice. For diagram planning, we experiment with three open-source LLMs: 1) LLama3 Huang et al. (2024), 2) Phi-3 Abdin et al. (2024) Mistral-7b Jiang et al. (2023) and for diagram rendering we use GPT4-o OpenAI (2023).
Critic Models for SciDoc2DiagrammarMAF.
Inspired by the success of GPT4 Vision in making pairwise comparison of evaluation of images using Question-Answering driven approaches OpenAI (2023); Zhang et al. (2023) and image aesthetic evaluation Abe et al. (2024), we adopt GPT4 Vision model to provide feedback on different aspects of our generated diagrams on various criteria (Prompts in  20,  21,  22).
@@
We have compared our approaches with existing text-to-image generation methods like DALLE3 Rombach et al. (2021) and Automatikz Belouadi et al. (2023). For experimenting with Diagram Planning in SciDoc2Diagrammer, we evaluate using zero-shot and few-shot (with 3 in-context exemplars) versions of LMs as mentioned in Table 8 (Prompts in 15, 16, 18, 19). We use Self-Refine by Madaan et al. (2023) as another baseline (Prompt in 25).
Figure 3: Average Human Rating on a scale of 1-5 on Completeness, Faithfulness and Layout on three subparts: SciDoc2DiagramBench-Gold (left), SciDoc2DiagramBench-Extended (middle) and SciMultiDoc2DiagramBench-Gold (right), it implies as the complexity of diagram creation increases, our proposed Multi-Aspect Refinement strategy appears to become more effective.
@@
However, simple token evaluations and neural evaluations like BERTScore and CLIPScore are only measuring string similarity and lacking a more finegrained evaluation framework that correlates more with human preferences Li et al. (2024c). Thus, we go beyond string similarity scores and include a comprehensive analysis of more finegrained aspects through human evaluation and the GPT4-V Evaluation. We access completeness (refers to the degree to which all relevant and necessary information is present in the generated figure), faithfulness (assesses how well the figure adheres to the facts, data, or specific instructions provided, ensuring that the content is correct and not misleading or misrepresented), and layout (refers to the visual clarity, focusing on how well the elements are structured and arranged) on a scale from 1 to 5.
@@ 6 Results and Findings
Our main objective is to explore the creation of ‘Extrapolated’ scientific visuals from documents using zero-shot/few-shot settings of advanced VLMs, and also analyze whether the quality of these visuals without dataset-specific fine-tuning, can be improved utilizing our proposed SciDoc2DiagrammarMAF. To answer this broad question, we focus our experiments on answering the following research questions as follows:
@@ 6.1 Main Research Questions and Results
GPT4-o is clearly the best-performing base/refiner model in terms of human and automatic judgement. In Table 13, GPT-4o consistently outperforms other models, including DALLE-3 and Automatikz, across all categories of diagrams on the SciDoc2DiagramBench. This superior performance is indicated by the green cells, which show GPT-4o achieving the highest scores in all the automatic metrics. We observe similar trend when we ask annotators to rate on the completeness, faithfulness and layout-satisfaction on 30 randomly selected samples of images from SciDoc2DiagramBench-Gold generated by each of the model (Figure 17).
RQ2: How effective is our SciDoc2Diagrammar-MAF over SciDoc2Diagrammer?
Based on Table 2, the effectiveness of our refinement strategies on SciDoc2DiagramBench-Gold utilizing GPT4-o with SumMAF and SeqMAF enhancements shows significant improvements over the Zero-shot (ZS) and Few-shot (FS) versions on all the diagram types on automatic evaluation. This finding holds true for SciMultiDoc2DiagramBench-Gold as well, as evident from the results in Table 7.
Drawing inspiration from Ribeiro et al. (2020), we acknowledge that automatic metrics alone are insufficient to fully evaluate model performance, so we also perform human evaluation. For SciDoc2DiagramBench Gold and SciMultiDoc2DiagramBench-Gold, we had human raters evaluate the quality of diagrams produced by different models and compare these to actual human-created slides, considering the source paper and the diagram’s intended purpose for 30 randomly sampled diagrams. In the SciMultiDoc2DiagramBench-Extended, which lacks a gold standard image, only the paper, the intent of the diagram, and the model-generated images were presented for evaluation on 50 randomly sampled images. Two expert co-authors and two professional crowdworkers from Upwork assessed the quality of all images, using criteria such as Completeness, Faithfulness, and Layout from 1 to 5. Figure 3 illustrates that diagrams generated using our refinement strategies compared to Few-shot and Self-Refine strategies consistently receive the highest scores from humans on all dataset subsets.
@@
According to Table 2, for flowcharts, SeqMAF shows significant improvement with the highest scores in all metrics compared to both ZS and FS. SumMAF also improves over ZS and FS, but to a lesser extent than SeqMAF, indicating SeqMAF’s superior refinement capability in terms of both completeness and detail accuracy. For Results diagrams, both SeqMAF and SumMAF show improvements over non-refined versions, with SeqMAF slightly leading in overall performance except for ClipScore. However, in the Architecture and Summary diagrams, SumMAF occasionally outperforms SeqMAF, particularly in BERTScore, and matches in ClipScore, indicating its strength in contexts requiring synthesized summarization of content. Since summarization-based refinement offers an integrated approach where multiple facets of feedback are considered simultaneously, this might have been more effective for diagrams requiring balanced presentation of dense information.
On SciDoc2DiagramBench-Extended, we explore whether human judgment perceives a significant enhancement in diagram quality across all categories on complex intents. We notice in Table 3, that even though there is a significant improvement of quality post-refinement in terms of human judgement on Faithfulness and Completeness (Flowcharts and tables improve significantly), Layout-satisfaction is the least (which can be attributed to subjectivity and LLM’s inability to refine on specific criteria).
@@
Across three datasets of increasing complexity, the Multi-Aspect Refinement (MAF-Refine) strategy consistently outperforms the few-shot approach in diagram production (Figure  3). On the simplest dataset, SciDoc2DiagramBench-Gold, MAF-Refine significantly betters the few-shot approach, with improvements of 0.9 in completeness, 1.1 in faithfulness, and 0.8 in layout. As complexity increases in SciDoc2DiagramBench-Extended, MAF-Refine still performs strongly, showing improvements of 1.4 in completeness, 1.0 in faithfulness, and 0.5 in layout. Even in the most complex dataset, SciMultiDoc2DiagramBench-Extended, MAF-Refine leads with improvements of 1.3 in completeness, 0.8 in faithfulness, and 0.3 in layout. These results demonstrate that MAF-Refine significantly boosts diagram quality across datasets and maintains its benefits even as complexity rises.
@@ 6.2 Ablation Analysis
Table 9 presents the performance impact of different components in the SciDoc2Diagrammer model using several metrics: CLIPScore (CS), BERTScore (BS), Completeness (C), Faithfulness (F), and Layout (L). Removing the Question Generation/Answering (QG/QA) component results in significant performance drops across all metrics, highlighting its importance. The absence of the Summarization component also negatively affects the scores but to a lesser extent.
@@
We conduct ablation analysis (Table 4) with SciDoc2Diagrammer-SeqMAF using BertScore (BS), Rouge (R) and ClipScore (CS) and consider SciDoc2DiagrammerBench-Gold. In particular, we experimented with the removal of each refinement module on the SciDoc2DiagramBench-Gold dataset, and analyzed which component is useful for what kind of diagram generation. We observe that completeness removal shows a significant drop in both summary and flowchart metrics, emphasizing the importance of comprehensive overviews and detailed visual representations. Faithfulness removal leads to a significant performance drop in results and flowcharts, underlining the necessity of faithful representation for coherent diagram creation. But layout removal has minimal impact on diagram quality improvement.
@@ 6.3 Qualitative Analysis
Taking a closer look at the diagrams generated before and after refinement (we sample the refined image having the highest average score rated by humans), we show in the figures 4, 19, 20 where our algorithm improves. For such qualitative analysis, we consider all four types of extrapolated images. In terms of precision, the refined versions of the diagrams appear to show a higher level of detail and precision. For example, in the block diagrams and flowcharts, the GPT4-o with refinement diagrams exhibit clearer layouts and better alignment of elements compared to those produced by baselines. The labels and connections between components are more accurate, suggesting that the refinement process improves the faithfulness and clarity of the diagrams. In scenarios where specific architectural details or complex data are necessary (as depicted in the block diagrams), the refined versions are closer to the original intent of diagram. This is particularly evident where GPT4-o with refinement accurately captures the necessary components and their interactions, unlike the outputs from baselines, which sometimes miss critical details or present a cluttered visualization.
@@ 7 Related Work
Learning from Feedback. Madaan et al. (2023) first introduced ‘Self-Refine’ and then there had been multiple works exploring the use of natural language feedback to improve performance. Lee et al. (2024a) proposes a framework to reduce hallucination in question answering. Ki and Carpuat (2024) exploits the strengths of LLMs and machine translation by guiding LLMs to automatically post-edit MT with external feedback on its quality. Lee et al. (2024b) introduced a multimodal self-feedback guided revision model to reduce hallucination. The work most closely related to ours is that of Nathani et al. (2023). However, our approach diverges from theirs in a key aspect: our feedback modules are uniquely grounded in cross-verifying information from long texts. Furthermore, we introduce a novel algorithm that refines this process sequentially, which has proven to be more effective in our specific use case. We also discuss the related works on code generation in 18.
@@ 8 Discussion and Future Work
We introduce an innovative task for constructing scientific diagrams from scientific document discourse, commonly used in presentations/posters. We establish a new benchmark using document-to-slides dataset and a sequential diagram refinement-based methodology to generate diagrams from academic papers. In future, we would like to assess the granularity of feedback that helps in diagram improvement by designing rubrics, framework, and guidelines about what a good diagram should look like by communicating with the experts in the data visualization communities. If we want to adopt data from a community, we also need to respect and adopt their evaluation rules and frameworks Li et al. (2024c). In addition, human-computer interaction is a hot topic as artificial intelligence (ai) advances in diverse tasks. However, ai still faces challenges such as gender and race discrimination Nghiem et al. (2024); An et al. (2024), producing unfaithful information through hallucinations Huang et al. (2023), privacy breaches Yao et al. (2024), etc. Fully relying on llms for diagram generation may overlook specific user intentions and use cases. Moreover, users learn their mental models of the data and refine their criteria and results through the interaction with ai Shankar et al. (2024); Li et al. (2024b). Future work could focus on interactive diagram generation, allowing users to iteratively engage with ai to refine and generate diagrams that best suit their specific use cases.
@@ Acknowledgement
We thank the anonymous ARR reviewers and UMD CLIP members for their constructive comments and feedback on the draft. This work was supported by a gift from Adobe Research.
@@
    Feedback Mechanism Limitations. While the multi-aspect feedback mechanism aims to enhance diagram quality by focusing on completeness, faithfulness, and layout, the effectiveness of this feedback is contingent upon the precision of the underlying evaluative criteria. Misalignments in feedback interpretation can lead to overfitting on specific diagram aspects at the expense of others.
@@
The experiments performed in this study involved human participants. All the experiments involving human evaluation in this paper were exempt under institutional IRB review. We recruited participants for our human study using Upwork and we have fairly compensated all the Upwork freelancers involved in this study, at an average rate of $20.00$ USD per hour (respecting their suggested Upwork hourly wage). We did not collect any personal data during the experiments, and they could choose not to participate in the study. The documents used in the study are distributed under an open license.

### v2

@@ 5 Experimental Setup
Flowchart Results Architecture Summary
@@
Model Flowchart Summary Architecture Results
@@ C F L C F L C F L C F L
For generating first-draft of diagrams, we experiment with state-of-the-art models to extract table from existing figures during diagram planning (Figure 2): 1) MatCha Liu et al. (2023b), 2) Chart-to-Text Kantharaj et al. (2022), 3) UniChart Masry et al. (2023), 4) DePlot Liu et al. (2023a) and GPT-4 Vision (GPT-4V) OpenAI (2023). Finally, based on our assessment (Figure 20), we concluded that GPT-4V outperformed other models and was the best choice. For diagram planning, we experiment with three open-source LLMs: 1) LLama3 Huang et al. (2024), 2) Phi-3 Abdin et al. (2024) Mistral-7b Jiang et al. (2023) and for diagram rendering we use GPT4-o OpenAI (2023).
Critic Models for SciDoc2Diagrammer-MAF.
Inspired by the success of GPT-4V in making pairwise comparison of evaluation of images using Question-Answering driven approaches OpenAI (2023); Zhang et al. (2023) and image aesthetic evaluation Abe et al. (2024), we adopt GPT-4V model to provide feedback on different aspects of our generated diagrams on various criteria (Prompts can be found in 20,  21,  22).
@@
We have compared our approaches with existing text-to-image generation methods like DALLE3 Rombach et al. (2021) and Automatikz Belouadi et al. (2023). For experimenting with Diagram Planning in SciDoc2Diagrammer, we evaluate using zero-shot and few-shot (with 3 in-context exemplars) versions of LMs as mentioned in Table 8 (Prompts in 15, 16, 18, 19). We use Self-Refine by Madaan et al. (2023) as another baseline (Prompt can be found in 25).
Figure 4: Average Human Rating on Completeness, Faithfulness and Layout on three subparts: SciDoc2DiagramBench-Gold (left), SciDoc2DiagramBench-Extended (middle) and SciMultiDoc2DiagramBench-Gold (right), it implies as the complexity of diagram creation increases, our proposed Multi-Aspect Refinement strategy appears to become more effective.
@@
However, simple token evaluations and neural evaluations like BERTScore and CLIPScore are only measuring string similarity and lacking a more finegrained evaluation framework that correlates more with human preferences Li et al. (2024c). Thus, we go beyond string similarity scores and include a comprehensive analysis of more finegrained aspects through human evaluation and the GPT4-V Evaluation. We assess completeness (measures the degree to which all relevant and necessary information is in the generated figure), faithfulness (assesses how well the figure adheres to the facts, data, or specific instructions provided, ensuring that the content is correct and not misleading or misrepresented), and layout (measures visual clarity, focusing on how well the elements are structured and arranged) on a scale from 1 to 5.
@@ 6 Results and Findings
Our main objective is to explore the creation of ‘Extrapolated’ scientific visuals from documents using zero-shot/few-shot settings of advanced VLMs, and also analyze whether the quality of these visuals without dataset-specific fine-tuning, can be improved using our proposed SciDoc2Diagrammer-MAF. To answer this broad question, we focus our experiments on answering the following research questions as follows:
@@ 6.1 Main Research Questions and Results
GPT4-o is clearly the best-performing base/refiner model on human and automatic judgement. In Table 13, GPT-4o consistently outperforms other models across all categories of diagrams on the SciDoc2DiagramBench. This superior performance is indicated by the green cells, which show GPT-4o achieving the highest scores in all the automatic metrics. We observe similar trend when we ask annotators to rate on the completeness, faithfulness and layout-satisfaction on 30 randomly selected samples of images from SciDoc2DiagramBench-Gold generated by each model (Figure 19).
RQ2: How effective is our Scidoc2diagrammer-MAF over SciDoc2Diagrammer?
Based on Table 2, the effectiveness of our refinement strategies on SciDoc2DiagramBench-Gold using GPT4-o with SumMAF and SeqMAF enhancements shows significant improvements over the Zero-shot (ZS) and Few-shot (FS) versions on all the diagram types on automatic evaluation. This finding holds true for SciMultiDoc2DiagramBench-Gold as well, as evident from the results in Table 7.
Drawing inspiration from Ribeiro et al. (2020), we acknowledge that automatic metrics alone are insufficient to fully evaluate model performance, so we also conduct human evaluation. For SciDoc2DiagramBench Gold and SciMultiDoc2DiagramBench-Gold, we had human raters evaluate the quality of diagrams produced by different models and compare these to actual human-created slides, considering the source paper and the diagram’s intended purpose for 30 randomly sampled diagrams. In the SciMultiDoc2DiagramBench-Extended, which lacks a gold standard image, only the paper, the intent of the diagram, and the model-generated images were presented for evaluation on 50 randomly sampled images. Two authors and two professional crowdworkers from Upwork assessed the quality of all images, using criteria such as Completeness, Faithfulness, and Layout from 1 to 5. Figure 4 illustrates that diagrams generated using our refinement strategies compared to Few-shot and Self-Refine strategies consistently receive the highest scores from humans on all dataset subsets.
Figure 5: Refinement of a flowchart for detecting rumors from Ma et al. (2017), highlighting improvements in clarity and completeness after refinement but emphasizing the continued absence of a critical feedback loop.
@@
From Table 2, SeqMAF shows significant improvement with the highest scores in all metrics compared to both ZS and FS for flowcharts. SumMAF also improves over ZS and FS, but to a lesser extent than SeqMAF, indicating SeqMAF’s superior refinement capability on both completeness and detail accuracy. For Results diagrams, both SeqMAF and SumMAF show improvements over non-refined versions, with SeqMAF slightly leading in overall performance except for ClipScore. However, in Architecture and Summary diagrams, SumMAF occasionally outperforms SeqMAF, indicating its strength in contexts requiring synthesized summarization of content. Since summarization-based refinement offers an integrated approach where multiple facets of feedback are considered simultaneously, this might have been more effective for diagrams requiring balanced presentation of dense information. On SciDoc2DiagramBench-Extended, we explore whether human judgment perceives a significant enhancement in diagram quality across all categories on complex intents. We notice in Table 3, that even though there is a significant improvement of quality post-refinement on human judgement on Faithfulness and Completeness (Flowcharts and tables improve significantly), Layout-satisfaction is the least (which can be attributed to subjectivity and LLM’s inability to refine on specific criteria).
@@
Across three datasets, the Multi-Aspect Refinement strategy consistently outperforms the few-shot approach in diagram production (Figure 4). On the simplest dataset, SciDoc2DiagramBench-Gold, MAF-Refine performs significantly better than few-shot approach, with improvements of 0.9 in completeness, 1.1 in faithfulness, and 0.8 in layout. As complexity increases in SciDoc2DiagramBench-Extended, MAF-Refine still performs strongly, showing improvements of 1.4 in completeness, 1.0 in faithfulness, and 0.5 in layout. Even in the most complex dataset, SciMultiDoc2DiagramBench-Extended, MAF-Refine leads with significant improvements, showing that MAF-Refine significantly boosts diagram quality across datasets and maintains its benefits even as complexity rises.
@@ 6.2 Ablation Analysis
Table 9 presents the performance impact of different components in the SciDoc2Diagrammer model using different metrics. Removing the Question Generation/Answering (QG/QA) component results in significant performance drops across all metrics, highlighting its importance. The absence of the Summarization component also negatively affects the scores but to a lesser extent.
@@
We conduct ablation analysis (Table 4) with SciDoc2Diagrammer-SeqMAF. In particular, we experimented with the removal of each refinement module on the SciDoc2DiagramBench-Gold dataset, and analyzed which component is useful to improve the diagram generation. Completeness removal shows a significant drop in both summary and flowchart metrics, emphasizing the importance of comprehensive overviews and detailed visual representations. Faithfulness removal leads to a significant performance drop in results and flowcharts, underlining the necessity of faithful representation for coherent diagram creation. But layout removal has the least impact on diagram quality improvement.
@@ 6.3 Qualitative Analysis
Qualitative analysis in feedback-based refinement (Figure 5) identifies gaps in the initial design, such as missing components or unclear integration of elements. After refinement, while the layout and completeness may improve significantly, issues like the absence of critical feedback loops can persist, leading to a less effective refinement process. This shows that even with visible improvements, missing critical features can detract from the model’s intended dynamic refinement functionality.
Besides, taking a closer look at the diagrams generated before and after refinement (we sample the refined image having the highest average score rated by humans), we show in Figure  6,  21,  22 where our algorithm improves. For such qualitative analysis, we consider all four types of extrapolated images. The refined versions of the diagrams appear to show a higher level of detail and precision. For example, in the block diagrams and flowcharts, the GPT4-o with refinement diagrams show clearer layouts and better alignment of elements compared to those produced by baselines. The labels and connections between components are more accurate, suggesting that the refinement process improves the faithfulness and clarity of the diagrams. In scenarios where specific architectural details or complex data are necessary, the refined versions are closer to the original intent of diagram. This is particularly evident where GPT4-o with refinement accurately captures the necessary components and their interactions, unlike the outputs from baselines, which sometimes present a cluttered visualization.
@@ 7 Related Work
Learning from Feedback. Madaan et al. (2023) first introduced ‘Self-Refine’ and then there had been multiple works exploring the use of natural language feedback to improve performance. Lee et al. (2024a) proposes a framework to reduce hallucination in question answering. Ki and Carpuat (2024) exploits the strengths of LLMs and machine translation by guiding LLMs to automatically post-edit MT with external feedback on its quality. Lee et al. (2024b) introduced a multimodal self-feedback guided revision model to reduce hallucination. The work most closely related to ours is that of Nathani et al. (2023). However, our approach diverges from theirs in a key aspect: our feedback modules are uniquely grounded in cross-verifying information from long texts. Furthermore, we introduce a novel algorithm that refines this process sequentially, which has proven to be more effective in our specific use case. We also discuss the related works on code generation in Appendix 18.
@@ 8 Discussion and Future Work
We introduce an innovative task for constructing scientific diagrams from scientific document discourse, commonly used in presentations/posters. We establish a new benchmark using document-to-slides dataset and a sequential diagram refinement-based methodology to generate diagrams from academic papers. In future, we would like to assess the granularity of feedback that helps in diagram improvement by designing rubrics about what a good diagram should look like by communicating with the experts in the data visualization communities.
@@ Acknowledgement
We thank the anonymous ARR reviewers and UMD CLIP members for their constructive comments and feedback on the draft. This material is based upon work supported by the National Science Foundation under Grant No. iis-2403436 (Boyd-Graber). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Annotations and support for Ishani Mondal were provided by a gift from Adobe Research.
@@
    Feedback Mechanism Limitations. While the multi-aspect feedback mechanism aims to enhance diagram quality by focusing on completeness, faithfulness, and layout, the effectiveness of this feedback depends on the precision of the underlying evaluative criteria. Misalignments in feedback interpretation can lead to overfitting on specific diagram aspects at the expense of others.
@@
The experiments performed in this study involved human participants. All the experiments involving human evaluation were exempt under institutional IRB review. We recruited participants for our human study using Upwork and we have fairly compensated all the Upwork freelancers involved in this study, at an average rate of $20.00$ USD per hour (respecting their suggested Upwork hourly wage). We did not collect any personal data during the experiments, and they could choose not to participate in the study. The documents used in the study are distributed under an open license.